One kilogram of oil has the following composition 45% carbon , 10% hydrogen , 15% sulfur , 15% methane , 6% carbon dioxide and 9% Nitrogen. Given that 25% of excess air is used for complete combustion. calculate the mass of the combustion products ( calculate only the mass of CO2 and SO2)
1
Expert's answer
2020-06-19T07:01:26-0400

Solution.

m(C) = 45 kg

m(S) = 15 kg

m(CH4) = 15 kg

m(CO2) = 6 kg

1) C + O2 = CO2

"n(C) = n(CO2) = 3746.7 \ mol"

m(CO2) = 165.0 kg

2) S + O2 = SO2

"n(S) = n(SO2) = 467.8 \ mol"

m(SO2) = 29.97 kg

3) CH4 + 2O2 = CO2 + 2H2O

"n(CH4) = n(CO2)"

m(CO2) = 41.12 kg

Sum(CO2) = 165.0 + 41.12 + 6.0 = 212.12 kg

Sum(SO2) = 29.97 kg

Answer:

Sum(CO2) = 212.12 kg

Sum(SO2) = 29.97 kg